LBTO Technical/D-Night

OSA: Steve Allanson

Isp: Jenny Power

Manager: John Little

ISA: B. Rothberg (Tucson - or an undisclosed location several miles below the surface)

================================================================

UT 12:23 - Configured for LBCs

UT 01:24 - Enclosure open

UT 01:51 - Going to SA113 to run DOFPIA

UT 02:02 - Rerunnig with X2 to get more pupils

UT 02:06 - Running co-pointing

UT 02:14 - Copointing done - having trouble with DS9 stretch for the LBC-Blue image for co-pointing

UT 02:16 - Running DOFPIA again - blue collimation iffy, at Airmass 1.43

UT 02:20 - Converged to 1" in LBC-R and 1.47" in Blue

UT 02:21 - Running RB_rV_sf7 PA=0

UT 02:27 - Running RB_rV_sf7_pa180

UT 02:35 - Running DOFPIA - can see wisps in images, may need to stop focus tests

UT 02:41 - looking dire with clouds - will run another SUPERFOC and see how it goes

UT 02:41 - Running RB_rV_sf7 PA=0

UT 02:49 - Running RB_rV_sf7_pa180

UT 02:50 - Running DOFPIA

UT 03:01 - Running RB_rv_sf7 PA=0

UT 03:07 - Running RB_rv_sf7_pa180

UT 03:14 - Running DOFPIA

UT 03:19 - Running RB_rv_sf7_pa90

UT 03:27 - Running RB_rv_sf7_pa270

UT 03:34 - Running DOFPIA

UT 03:39 - Running RB_rv_sf7_pa90

Running RB_rv_sf7_pa270

Running DOFPIA

UT 03:59 - Running RB_rv_sf7_pa90

UT 04:06 - Running RB_rv_sf7_270

UT 04:11 - Counts higher, but clouds increasing rapidly last 10min - could be the end of LBC testing for now, will complete final set of foc

UT 04:18 - Switching to MODS1 due to wispy clouds and scattered light from bright Moon

UT 04:23 - Waking MODS1

UT 04:42 - Collimating -Guider shows avg 1.7", WFS 1.4" and DIMM 1.7"

===========================================================

UT 04:43 - Starting acq IR20087-0308 ULIRG - Rothberg

UT 04:55 - Starting science script - Airmass 1.4 - seeing 2"

UT 04:58 - Clouds have cleared off most of sky, but seeing is awful

UT 06:00 - Doing 1 more 1800sec exposure

UT 06:14 - Seeing blown up to 2" avg guider 1.8" WFS

UT 06:33 - Moving to Feige 110 SpecStd

UT 06:39 - Started Feige 110 spectro script

UT 06:51 - Moving to FF0234-0139 Rothberg mid-z ULIRG

UT 06:53 - Airmass 1.530

UT 07:02 - in two minutes cannot see target in acquisition image

UT 07:03 - Taking a 3minute acq to see if smudge is real

UT 07:08 - Think I see the smudge in 3min

UT 07:09 - guider, WFS all jumping around from 1.3-2.5"

UT 07:14 - Giving up - seeing too unstable going to backup

UT 07:15 - Moving to UGC 2238 - going to finish what I started on Sep 25

UT 07:26 - Starting science

UT 08:09 - Seeing seems to have calmed down a bit, once UGC2238 finishes, will switch to LBCs

UT 08:32 - Completed - switching back to MODS1

UT 08:33 - Putting MODS1 to sleep

====================================================================

UT 08:34 - Reconfiguring to LBC

UT 08:46 - Going to Standard field SA95 to do focus

UT 08:48 - Running DOFPIA
UT 08:55 - Red to 0.97" Blue to 1.14"

UT 08:55 - Running through all filters in SA95 Airmass 1.323

UT 09:01 - Hardware Failure - timeout error on Filter LBC-R going from z-sloan to F972N

UT 09:03 - Pressed "Play" and it continued OK

=====================================================================

UT 09:20 - Going to M31substr3 (LBTB field) - made a shorter set of exposures, each dither has:

60x10 and 30x4 for V;

150x2 for TiO and CN

35 x3 for I

-same total exposure time as original script

UT 09:23 - Running DOFPIA

UT 09:27 - LBC-B converged to 1.12", LBC-R 1"

UT 09:29 - Starting M31substr3_VshorterV2

UT 09:32 - getting ~20k background V-band 60sec, 100sec TiO 7k

UT 09:34 - DIMM jumped, when started was 1" now saying 1.4"

UT 09:42 - Guider LBC-R reporting ~1" again

UT 09:45 - CN ~5k counts bkgd, I-band in 35 sec ~14k counts

UT 09:50 - Blue CCD Controller error after 21/69 exposures

2015/09/29 09:50:14.675251 N B completed 21 OB image/s
2015/09/29 09:50:14.675286 E B OB execution error: one or more systems failed [src/lbc/actions.c:1160]
2015/09/29 09:50:14.675400 N - freeing FITS headers
2015/09/29 09:50:14.675724 N - UI update with message >(BLUE Channel) OB execution error: (CAMERA) CCD Controller error!< at 1443520214.675504
2015/09/29 09:50:56.131000 E B CAMERA CCDCTRL "FPGA not loaded" (code:48) controller error [ccdctrl.c:2439]
UT 09:51 - Turning off LBC

UT 09:52 - Turning ON LBC

UT 09:57 - Taking opportunity to DOFPIA

UT 09:58 - LBC crash was 2 exposures shy of completing first dither position. Going to continue from dither position 2 and do each dither separately (already made OBs for them) and DOFPIA in between

UT 10:05 - starting M31substr3_Vshorter_dither2

UT 10:07 - Red reporting 2" lbcr.20150929.100539 - but imexam on actual image shows ~0.8"

UT 10:17 - LBC -B issue - stopped at 6/10 in V (finished 8/23 total)

2015/09/29 10:17:43.614000 CAMERA CCDCTRL "FPGA not loaded" (code:48) controller error [ccdctrl.c:2439]
78 2015/09/29 10:16:57.333836 OB execution error: one or more systems failed [src/lbc/actions.c:1160]
77 2015/09/29 10:16:57.232784 >>> preset/expose/save interrupted at preset time, retcode=7
76 2015/09/29 10:16:57.232663 camera system encountered a problem 9 [src/channel/channel.c:1132]
75 2015/09/29 10:16:57.232573 CAMERA CCD Controller error (8) [src/camera/camera.c:1146]
74 2015/09/29 10:16:57.232529 CAMERA reinitializing CCD Controller failed
73 2015/09/29 10:16:58.679000 CAMERA CCDCTRL load FPGA on SPC-A200 error (timeout reading SPC status) [ccdctrl.c:934]

IT #2397

UT 10:19 - Stopping LBC and turning off

UT 10:24 - lbcb.201509.101503 shows striping

UT 10:25 - Running DOFPIA

UT 10:30 - ReRunning M31substr3_Vshorter_dither2

UT 10:42 - Seeing is blowing up during exposures-Blue seems better according to guider - going to continue at least until end of OB - however actual images seem about 1-1.3"

UT 10:54 - Running DOFPIA to gauge conditions

UT 11:01 - Going back to SA95

UT 11:22 - Taking long exposures with CN and TiO to get stars on Tekchips

UT 11:40 - Did 3/4 dither positions on SA95 - should be enough to get a bunch of TEK chips with guider stars in known field

UT 11:40 - Moving back to 3rd dither position on M31substr3

UT 11:42 - Running DOFPIA

UT 11:46 - Running M31substr3_Vshorter_dither3 - trying to take advantage of late good seeing DIMM 1"

=======================================================

UT 12:09 - Attempting Veillet recovery BZ509

UT 12:29 - We had another LBC filter wheel issue - 10/20 completed, however on recovery I didn't hit the NSIGUI hijack so last 4 frames are saturated and/not non-sidereal guiding

UT 12:30 - DONE

UT 12:32 - disconnecting and powering off LBC

-- %USERSIG{BarryRothberg - 2015-09-29}%

Comments

 
Topic revision: r6 - 29 Sep 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback